# Consumption Partial Insurance

### v5

@@ B.2 Linear consumption function parameters
Given the income process, the linear consumption function, and the properties of shocks, the transmission parameters are given by
$\phi^{(1)}_{t}=\frac{1}{2}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\Delta^{2}y_{it+\kappa})/\sigma_{\zeta_{t}}^{2}\qquad\text{and}\qquad\psi^{(1)}_{t}=-{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it+2})/\sigma_{v_{t}}^{2}.$
@@
$\sigma_{\xi_{t}}^{2}=\frac{1}{2}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\Delta^{2}c_{it+\kappa})-(\phi^{(1)}_{t})^{2}\sigma_{\zeta_{t}}^{2}-\frac{1}{2}(\psi^{(1)}_{t})^{2}\sigma_{v_{t}}^{2}\quad\text{and}\quad\sigma_{u^{c}_{t}}^{2}=-{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it+2}).$
@@ B.3 Quadratic consumption function parameters
Given the income process, the quadratic consumption function, and the properties of income shocks, the transmission parameters are given by
@@
$\left(\begin{array}[]{l}\phi^{(1)}_{t}\\ \psi^{(1)}_{t}\\ \phi^{(2)}_{t}\\ \psi^{(2)}_{t}\\ \omega^{(22)}_{t}\end{array}\right)=\mathbf{A}^{-1}\left(\begin{array}[]{l}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it+2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it+2})^{2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it}\times\Delta^{2}y_{it+2})\\ \end{array}\right),$ (B.1)
@@
where the matrix of coefficients $\mathbf{A}$ depends exclusively on the second, third, and fourth moments of income shocks. [^1] With biennial data, matrix $\mathbf{A}$ is different from $\mathbf{A}$ with annual data in the main paper and is given by $\mathbf{A}$ $=\left(\begin{array}[]{lllll}2\sigma_{\zeta_{t}}^{2}&\sigma_{v_{t}}^{2}&2\gamma_{\zeta_{t}}&\gamma_{v_{t}}&0\\ 2\gamma_{\zeta_{t}}&\gamma_{v_{t}}&2(\kappa_{\zeta_{t}}+\sigma_{\zeta_{t}}^{2}(3\sigma_{\zeta_{t}}^{2}+2\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2}))&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(2\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2})&4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ 0&-\sigma_{v_{t}}^{2}&0&-\gamma_{v_{t}}&0\\ 0&\gamma_{v_{t}}&4\sigma_{\zeta_{t}}^{2}(\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(2\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2})&0\\ 0&-\gamma_{v_{t}}&-2\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2})&-\kappa_{v_{t}}-\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}&-2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ \end{array}\right).$ The determinant of this matrix is given by
$\text{det}(\mathbf{A})$ $=$ $-8\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}\Big{[}\Big{(}k_{\zeta_{t}}$ $-\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+5\sigma_{\zeta_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}\Big{(}k_{v_{t}}-\frac{\gamma_{v_{t}}^{2}}{\sigma_{v_{t}}^{2}}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+\sigma_{v_{t}}^{4}+2\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}$
$-8\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{6}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{4}-8\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{4}\sigma_{u^{y}_{t}}^{2}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{]}.$
@@
Assuming that the shocks are not drawn from a Bernoulli distribution, kurtosis is bounded below by $\gamma^{2}/\sigma^{2}+\sigma^{4}$. It follows that
$k_{\zeta_{t}}-\gamma_{\zeta_{t}}^{2}/\sigma_{\zeta_{t}}^{2}-\sigma_{\zeta_{t}}^{4}$ $>0,$ (B.2)
$k_{v_{t}}-\gamma_{v_{t}}^{2}/\sigma_{v_{t}}^{2}-\sigma_{v_{t}}^{4}$ $>0.$ (B.3)
@@
$6\sigma_{\zeta_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+2\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $>0,$ (B.4)
$2\sigma_{v_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+2\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $>0.$ (B.5)
Adding the inequalities (B.2)+(B.4) and (B.3)+(B.5) by parts, it follows that $k_{\zeta_{t}}-\gamma_{\zeta_{t}}^{2}/\sigma_{\zeta_{t}}^{2}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+5\sigma_{\zeta_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}>0$ and $k_{v_{t}}-\gamma_{v_{t}}^{2}/\sigma_{v_{t}}^{2}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+\sigma_{v_{t}}^{4}+2\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}>0$, so the bracketed terms in the first line of the determinant are both strictly positive. It then follows that
$\Big{(}k_{\zeta_{t}}-\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+5\sigma_{\zeta_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}\Big{(}k_{v_{t}}-\frac{\gamma_{v_{t}}^{2}}{\sigma_{v_{t}}^{2}}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+\sigma_{v_{t}}^{4}+2\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}$
$-8\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{6}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{4}-8\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{4}\sigma_{u^{y}_{t}}^{2}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $>$
$\Big{(}\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}+\sigma_{\zeta_{t}}^{4}-\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+5\sigma_{\zeta_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}\Big{(}\frac{\gamma_{v_{t}}^{2}}{\sigma_{v_{t}}^{2}}+\sigma_{v_{t}}^{4}-\frac{\gamma_{v_{t}}^{2}}{\sigma_{v_{t}}^{2}}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+\sigma_{v_{t}}^{4}+2\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}$
$-8\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{6}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{4}-8\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{4}\sigma_{u^{y}_{t}}^{2}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $=$
$\Big{(}6\sigma_{\zeta_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+2\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}\Big{(}2\sigma_{v_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+2\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}$
$-8\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{6}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{4}-8\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{4}\sigma_{u^{y}_{t}}^{2}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $=$
$4\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}+4\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}+4\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}\sigma_{u^{y}_{t}}^{2}$ $>0,$
@@
where the first inequality follows from the lower bound on the kurtosis. The term inside the squared brackets in the determinant is thus strictly positive, so $\text{det}(\mathbf{A})<0$. The matrix of coefficients is thus full rank and invertible for any distribution of permanent & transitory shocks except Bernoulli. [^2] If the distributions of both permanent and transitory shocks were Bernoulli, kurtosis would equal its lower bound $\gamma^{2}/\sigma^{2}+\sigma^{4}$ and the determinant would not always be strictly away from zero.
@@
How do different levels of skewness and kurtosis influence the transmission parameters? Consider the transmission of permanent shocks, abstracting from transitory shocks, i.e., $\psi^{(1)}_{t}=\psi^{(2)}_{t}=\omega^{(22)}_{t}=0$. $\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ are pinned down by the first equation in (B.1), namely $2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}+2\gamma_{\zeta_{t}}\phi^{(2)}_{t}={\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})$. Absent any transmission of transitory shocks, the right hand side reflects the strength of the covariance between consumption growth and permanent shocks. Suppose $\phi^{(1)}_{t}$ is fixed. We highlight three cases:
@@
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}=0$. Letting $\gamma_{\zeta_{t}}<0$ as in the data, this equation restricts $\phi^{(2)}_{t}=0$. Tail shocks thus transmit similarly to non-tail ones, i.e., this configuration identifies a linear consumption process.
@@
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}>0$. Consumption growth and permanent shocks co-vary more strongly than the linear consumption process suggests, i.e., there is excess co-movement. For $\gamma_{\zeta_{t}}<0$, this configuration identifies $\phi^{(2)}_{t}<0$ and implies that negative shocks transmit more strongly than other shocks. Negative shocks are often the largest shocks (due to left skewness), which explains the excess co-movement. Ceteris paribus, a less skewed distribution ($\gamma_{\zeta_{t}}$ drops in magnitude) implies a more negative $\phi^{(2)}_{t}$, because a more symmetric distribution associated with fewer large shocks requires a stronger transmission of said shocks to maintain a given excess co-movement. If $\gamma_{\zeta_{t}}>0$, this configuration identifies $\phi^{(2)}_{t}>0$ and analogous arguments apply to positive shocks.
@@
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}<0$. Consumption growth and permanent shocks co-move less strongly than the linear process suggests, i.e., there is excess smoothness. For $\gamma_{\zeta_{t}}<0$, this identifies $\phi^{(2)}_{t}>0$ and suggests a muted pass-through of negative shocks. These are the largest shocks to income, so muting their pass-through suppresses the covariance between consumption and income growth. The opposite is true if $\gamma_{\zeta_{t}}>0$.

### v6

@@ B.2 Linear consumption function parameters
Given the general income process, the linear consumption function, and the properties of shocks, the transmission parameters are given by
$\phi^{(1)}_{t}$ $=\frac{1}{(1+\rho^{2})\sigma_{\zeta_{t}}^{2}}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\rho^{-\kappa}\Delta^{2}_{\rho}y_{it+\kappa}),$
$\psi^{(1)}_{t}$ $=-\frac{1}{\rho^{2}\sigma_{v_{t}}^{2}}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it+2}).$
@@
$\sigma_{\xi_{t}}^{2}$ $=\frac{1}{2}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\Delta^{2}c_{it+\kappa})-(\phi^{(1)}_{t})^{2}\frac{1+\rho^{2}}{2}\sigma_{\zeta_{t}}^{2}-\frac{1}{2}(\psi^{(1)}_{t})^{2}\sigma_{v_{t}}^{2},$
$\sigma_{u^{c}_{t}}^{2}$ $=-{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it+2}).$
@@ B.3 Quadratic consumption function parameters
Given the general income process, the quadratic consumption function, and the properties of shocks, the transmission parameters are given by
@@
$\left(\begin{array}[]{l}\phi^{(1)}_{t}\\ \psi^{(1)}_{t}\\ \phi^{(2)}_{t}\\ \psi^{(2)}_{t}\\ \omega^{(22)}_{t}\end{array}\right)=\mathbf{A}^{-1}\left(\begin{array}[]{l}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}_{\rho}y_{it})^{2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it+2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}_{\rho}y_{it+2})^{2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2})\\ \end{array}\right),$ (B.1)
@@
where the matrix of coefficients $\mathbf{A}$ depends exclusively on the second, third, and fourth moments of income shocks, the AR(1) persistence parameter $\rho$, and the variance of income measurement error. [^1] While the paper shows matrix $\mathbf{A}$ for the case of annual data, unit root in income, and no measurement error, the general form of $\mathbf{A}=[a_{ij}]_{5\times 5}$ with biennial data, income persistence, and measurement error has $a_{15}=a_{31}=a_{33}=a_{35}=a_{41}=a_{45}=a_{51}=0$, and $a_{11}$ $=(1+\rho^{2})\sigma_{\zeta_{t}}^{2},$ $a_{12}$ $=\sigma_{v_{t}}^{2},$ $a_{13}$ $=(1+\rho^{3})\gamma_{\zeta_{t}},$ $a_{14}$ $=\gamma_{v_{t}},$ $a_{21}$ $=(1+\rho^{3})\gamma_{\zeta_{t}},$ $a_{22}$ $=\gamma_{v_{t}},$ $a_{23}$ $=(1+\rho^{4})\kappa_{\zeta_{t}}+6\rho^{2}(\sigma_{\zeta_{t}}^{2})^{2}+(1+\rho^{2})(1+\rho^{4})\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2}),$ $a_{24}$ $=\kappa_{v_{t}}+\rho^{4}(\sigma_{v_{t}}^{2})^{2}+(1+\rho^{2})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+(1+\rho^{4})\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2},$ $a_{25}$ $=2(1+\rho^{2})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2},$ $a_{32}$ $=-\rho^{2}\sigma_{v_{t}}^{2},$ $a_{34}$ $=-\rho^{2}\gamma_{v_{t}},$ $a_{42}$ $=\rho^{4}\gamma_{v_{t}},$ $a_{43}$ $=(1+\rho^{2})^{2}(\sigma_{\zeta_{t}}^{2})^{2}+(1+\rho^{2})(1+\rho^{4})\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2}),$ $a_{44}$ $=\rho^{4}\kappa_{v_{t}}+(\sigma_{v_{t}}^{2})^{2}+(1+\rho^{2})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+(1+\rho^{4})\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2},$ $a_{52}$ $=-\rho^{2}\gamma_{v_{t}},$ $a_{53}$ $=-\rho^{2}(1+\rho^{2})\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2}),$ $a_{54}$ $=-\rho^{2}(\kappa_{v_{t}}+\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}),$ $a_{55}$ $=-\rho^{2}(1+\rho^{2})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}.$ Its determinant [^2] The simpler derivation for $\rho=1$ appears in older versions of the paper, see arxiv.org/abs/2306.13208. is given by
$\text{det}(\mathbf{A})$ $=-\rho^{4}(1+\rho^{2})\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}\times$
$\bigg\{$ $\bigg((1+\rho^{2})(1+\rho^{4})(k_{\zeta_{t}}-\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}-\sigma_{\zeta_{t}}^{4})+(\rho^{2}-\rho)^{2}\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}$
$+\frac{1+\rho^{2}}{\rho^{4}}(1+2\rho^{2}+2\rho^{4}+6\rho^{6}+\rho^{8})\sigma_{\zeta_{t}}^{4}+\frac{(1+\rho^{2})^{2}}{\rho^{4}}(1+\rho^{8})\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2})\bigg)\times$
$\bigg(\rho^{4}(k_{v_{t}}-\frac{\gamma_{v_{t}}^{2}}{\sigma_{v_{t}}^{2}}-\sigma_{v_{t}}^{4})+(1+\rho^{4})\sigma_{v_{t}}^{4}+(1+\rho^{2})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+(1+\rho^{4})\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\bigg)$
$-\frac{(1+\rho^{2})^{4}}{\rho^{4}}(1+\rho^{4})\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}-\frac{(1+\rho^{2})^{2}}{\rho^{4}}(1+\rho^{4})(1+\rho^{8})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{6}$
$-2\frac{(1+\rho^{2})^{3}}{\rho^{4}}(1+\rho^{4}+\rho^{8})\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}-\frac{(1+\rho^{2})^{2}}{\rho^{4}}(1+\rho^{4})(1+\rho^{8})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{4}$
$-2\frac{(1+\rho^{2})^{2}}{\rho^{4}}(1+\rho^{4})(1+\rho^{8})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{4}\sigma_{u^{y}_{t}}^{2}-2\frac{(1+\rho^{2})^{3}}{\rho^{4}}(1+\rho^{4}+\rho^{8})\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\bigg\}.$
@@
Assuming that the shocks are not drawn from a Bernoulli distribution, kurtosis is bounded below by $\gamma^{2}/\sigma^{2}+\sigma^{4}$. Given that $0<\rho\leq 1$, it follows that
$(1+\rho^{2})(1+\rho^{4})(k_{\zeta_{t}}-\gamma_{\zeta_{t}}^{2}/\sigma_{\zeta_{t}}^{2}-\sigma_{\zeta_{t}}^{4})$ $>0,$ (B.2)
$\rho^{4}(k_{v_{t}}-\gamma_{v_{t}}^{2}/\sigma_{v_{t}}^{2}-\sigma_{v_{t}}^{4})$ $>0.$ (B.3)
@@
$(\rho^{2}-\rho)^{2}\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}$ $\geq 0,$ (B.4)
$\frac{1+\rho^{2}}{\rho^{4}}(1+2\rho^{2}+2\rho^{4}+6\rho^{6}+\rho^{8})\sigma_{\zeta_{t}}^{4}+\frac{(1+\rho^{2})^{2}}{\rho^{4}}(1+\rho^{8})\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2})$ $>0,$ (B.5)
$(1+\rho^{4})\sigma_{v_{t}}^{4}+(1+\rho^{2})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+(1+\rho^{4})\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $>0,$ (B.6)
as the sum of mostly strictly positive terms. Adding the inequalities (B.2)+(B.4)+(B.5) and (B.3)+(B.6) by parts, it follows that the two long bracketed terms in the first lines of the determinant are strictly positive, so their product is also strictly positive. Given the lower bound on the kurtosis, the term inside the curly brackets is then strictly greater than
$\bigg((\rho^{2}-\rho)^{2}\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}+\frac{1+\rho^{2}}{\rho^{4}}(1+2\rho^{2}+2\rho^{4}+6\rho^{6}+\rho^{8})\sigma_{\zeta_{t}}^{4}+\frac{(1+\rho^{2})^{2}}{\rho^{4}}(1+\rho^{8})\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2})\bigg)$ $\times$
$\bigg((1+\rho^{4})\sigma_{v_{t}}^{4}+(1+\rho^{2})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+(1+\rho^{4})\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\bigg)$
$-\frac{(1+\rho^{2})^{4}}{\rho^{4}}(1+\rho^{4})\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}-\frac{(1+\rho^{2})^{2}}{\rho^{4}}(1+\rho^{4})(1+\rho^{8})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{6}$
$-2\frac{(1+\rho^{2})^{3}}{\rho^{4}}(1+\rho^{4}+\rho^{8})\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}-\frac{(1+\rho^{2})^{2}}{\rho^{4}}(1+\rho^{4})(1+\rho^{8})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{4}$
$-2\frac{(1+\rho^{2})^{2}}{\rho^{4}}(1+\rho^{4})(1+\rho^{8})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{4}\sigma_{u^{y}_{t}}^{2}-2\frac{(1+\rho^{2})^{3}}{\rho^{4}}(1+\rho^{4}+\rho^{8})\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $=$
$(\rho^{2}-\rho)^{2}\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}\bigg((1+\rho^{4})\sigma_{v_{t}}^{4}+(1+\rho^{2})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+(1+\rho^{4})\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\bigg)$
$+4\rho^{2}(1+\rho^{2})^{2}\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}+4\rho^{2}(1+\rho^{2})(1+\rho^{4})\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}+4\rho^{2}(1+\rho^{2})(1+\rho^{4})\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $>0.$
@@
The last inequality follows from $0<\rho\leq 1$ and the fact that we are adding mostly strictly positive terms (and one weakly positive, the term involving $\gamma_{\zeta_{t}}^{2}$). The expression inside the curly brackets in the determinant is thus strictly positive, so $\text{det}(\mathbf{A})<0$. $\mathbf{A}$ is thus full rank and invertible for any distribution of permanent & transitory shocks except Bernoulli. [^3] If the distributions of both permanent and transitory shocks were Bernoulli, kurtosis would equal its lower bound $\gamma^{2}/\sigma^{2}+\sigma^{4}$ and the determinant may not always be strictly away from zero.
@@
How do different levels of skewness and kurtosis influence the transmission parameters? Consider the transmission of permanent shocks, abstracting from transitory shocks, i.e., $\psi^{(1)}_{t}=\psi^{(2)}_{t}=\omega^{(22)}_{t}=0$, and let $\rho=1$. $\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ are pinned down by the first equation in (B.1), namely $2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}+2\gamma_{\zeta_{t}}\phi^{(2)}_{t}={\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it})$. Absent transmission of transitory shocks, the right hand side reflects the strength of the covariance between consumption growth and permanent shocks. Suppose $\phi^{(1)}_{t}$ is fixed. We highlight three cases:
@@
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}=0$. Letting $\gamma_{\zeta_{t}}<0$ as in the data, this equation restricts $\phi^{(2)}_{t}=0$. Tail shocks thus transmit similarly to non-tail ones, i.e., this configuration identifies a linear consumption process.
@@
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}>0$. Consumption growth and permanent shocks co-vary more strongly than the linear consumption process suggests, i.e., there is excess co-movement. For $\gamma_{\zeta_{t}}<0$, this configuration identifies $\phi^{(2)}_{t}<0$ and implies that negative shocks transmit more strongly than other shocks. Negative shocks are often the largest shocks (due to left skewness), which explains the excess co-movement. Ceteris paribus, a less skewed distribution ($\gamma_{\zeta_{t}}$ drops in magnitude) implies a more negative $\phi^{(2)}_{t}$, because a more symmetric distribution associated with fewer large shocks requires a stronger transmission of said shocks to maintain a given excess co-movement. If $\gamma_{\zeta_{t}}>0$, this configuration identifies $\phi^{(2)}_{t}>0$ and analogous arguments apply to positive shocks.
@@
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}<0$. Consumption growth and permanent shocks co-move less strongly than the linear process suggests, i.e., there is excess smoothness. For $\gamma_{\zeta_{t}}<0$, this identifies $\phi^{(2)}_{t}>0$ and suggests a muted pass-through of negative shocks. These are the largest shocks to income, so muting their pass-through suppresses the covariance between consumption and income growth. The opposite is true if $\gamma_{\zeta_{t}}>0$.